Can anybody help me Please
Find a polynomial with zeros of 2-3i and -3+5i and multiply it out

Answer by venugopalramana(3286)

Can anybody help me Please
Find a polynomial with zeros of 2-3i and -3+5i and multiply it out
IF THE POLYNOMIAL HAS REAL COEFFICIENTS THEN 2+3I AND -3-5I ARE ALSO THE ZEROS
HENCE THE POLYNOMIAL IS
F(X) = [X-(2-3I)][X-(2+3I)][X-(-3+5I)][X-(-3-5I)]=0
F(X)=[(X-2)^2-(3I)^2][(X+3)^2-(5I)^2]=0
F(X)=[X^2-4X+4+9][X^2+6X+9+25]=0
F(X)=X^4+6X^3+34X^2-4X^3-24X^2-136X+13X^2+78X+442=0
F(X)=X^4+2X^323X^2-58X+442=0